A student throws a set of keys vertically upward to his flatemate, who is in a window 3,.75 m above. The flatemates outstretched hand catches the keys 1.70 s later. (Take upward as the positive direction. Indicate the direction with the sign of your answer.) O 012 points Previous Answers SerCP11GE 2.Po59 (a) With what initial velocity were the keys thrown? 179 Your response differs from the correct answer by more than 10%. Double check your calculations m/s (b) What was the velocity of the keys just before they were caught? -125 The response you submitted has the wrong sign, m/s eBock
Answer #1

Solution) Given y=3.75m

t=1.70 s

(a) initial velocity Vi = ?

We have y = Vi(t) + (1/2)gt^2

Since thrown upwards so g is negative

3.75 = Vi(1.70) + (0.5×(-9.8)×(1.70^2))

3.75 =(1.70)Vi - (4.9×1.70^2)

17.91 = (1.70)Vi

Vi = 10.53 m/s upward direction

(b) velocity just before they were caught V =?

V = Vi + gt

V= 10.53 + (-9.8)×(1.70)

V= - 6.13 m/s downward direction